OpenStudy (anonymous):

there are 5students in a group find the number of students in 5 groups

OpenStudy (anonymous):

there would be 25 students

OpenStudy (anonymous):

5 x 5 = 25

OpenStudy (anonymous):

5*5=25

OpenStudy (anonymous):

there are 70pages in a notebook find the number of pages on 8 notenooks

OpenStudy (anonymous):

8 x 70

OpenStudy (anonymous):

70*8 is 560 so there would be 560 pages in 8 notebooks

OpenStudy (anonymous):

find each unit price $5 for 10 pounds

OpenStudy (anonymous):

Any time you have a problem where you're told that there are a certain number of things in something and there are a number of those somethings, you want to do multiplication.

OpenStudy (anonymous):

If it's 5 for 10, then it's 5/10 for one.

OpenStudy (anonymous):

1 dollar for 2 pounds or 50 cents for 1 pound

OpenStudy (anonymous):

40 ounces for $12

OpenStudy (anonymous):

Here they reversed it to try to trick you! It's $12 for 40, so it's 12/40 for one. 12/40 = 6/20 = 3/10 or 30 cents.

OpenStudy (anonymous):

$6 for 10 pens

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Try it...look at the last two and make an intelligent guess.

OpenStudy (anonymous):

i thinks 25

OpenStudy (anonymous):

If $5 for 10 was 5/10, and $12 for 40 was 12/40, what do you think $6 for 10 is?

OpenStudy (anonymous):

6/10

OpenStudy (anonymous):

There you go! 6/10 = 60 cents.

OpenStudy (anonymous):

what about $60 for 5 books

OpenStudy (anonymous):

You tell me

OpenStudy (anonymous):

60/5=12

OpenStudy (anonymous):

Nice!

OpenStudy (anonymous):

and $22 for 3 shirts

OpenStudy (anonymous):

A little tricky this one. What do you think?

OpenStudy (anonymous):

22/3=7.3

OpenStudy (anonymous):

Wow, very impressive! Since it's in dollars an cents, it would actually be $7.33 but you're absolutely correct.

OpenStudy (anonymous):

Last one...I have to get back to work.

OpenStudy (anonymous):

25/100,4/16 do the ratios in each pair form a proportion

OpenStudy (anonymous):

Can you simplify 25/100?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Same as 4/16, right?

OpenStudy (anonymous):

nope its 1/4

OpenStudy (anonymous):

Correct. So is 4/16 if you simplify it.

OpenStudy (anonymous):

yes but i still need help

OpenStudy (anonymous):

You've got your ratio: 1/4.

OpenStudy (anonymous):

i know i have more homework to do

OpenStudy (anonymous):

They're both the same, so the proportion is 1, right?

OpenStudy (anonymous):

Aha...post it to the site...I know others will offer help. And give yourself credit for how intelligent you are. There are enough people in the world to tell you (or me) that we're not smart enough. Don't be one of them.

OpenStudy (anonymous):

Take care.

OpenStudy (anonymous):

15/20,4/5

OpenStudy (anonymous):

(Post in a new question or nobody will see it)